OpenStudy (anonymous):

whats the integral of xe^(4-x) as x approaches infinity?

OpenStudy (mathteacher1729):

\[\int_{}^{\infty} xe^{4-x}dx\] What is the lower limit? really can't do much without that.

OpenStudy (anonymous):

oops, sorry! 4

OpenStudy (mathteacher1729):

\[\int_{4}^{\infty} xe^{4-x}dx\] is that the problem?

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

its not infinity or -infinity, but I cannot find out the answer

OpenStudy (anonymous):

Integrate by parts u=x and dv=e^4-x change infinity to t, then evaluate as a limit

OpenStudy (anonymous):

(as t goes to infinity)

OpenStudy (anonymous):

\[\lim_{4 \rightarrow \infty} -x e^{4-x} -e^{4-x} ?\]

OpenStudy (anonymous):

? i have no idea...

OpenStudy (anonymous):

x--->infinity

OpenStudy (anonymous):

*

OpenStudy (anonymous):

Yep, you got the integral right. now sub in your upper and lower limits. Now you should have something with t in it ( y(t) ) Take the limit of this as t goes to infinity